Given the equation C(g)=3.03(g)...what does the number 3.03 represent?...Find C(2)...name one value that would be inappropriate for this function's purpose...graphing C(g), what would be the domain...range...find the slope:coordinates(1997, 1.26) and (2006, 2.31)If you can help me I would be forever grateful.
Answer by stanbon(75887)

Given the equation C(g)=3.03(g)...what does the number 3.03 represent?
Ans: 3.03 is the slope of the linear function.
-----------------------------------------------------------
Find C(2)
Ans: C(2) = 3.03*2 = 6.06
-------------------------------
Name one value that would be inappropriate for this function's purpose...graphing C(g)
Ans: There are no values of "g" that are inappropriate.
------------------------------
What would be the:
domain: All Real Numbers
range: All Real Numbers
---------------------------------
Find the slope:coordinates(1997, 1.26) and (2006, 2.31)
slope = [2.31 - 1.26]/[2006 - 1997] = 1.05/9 = 0.1166666...
============================
Cheers,
Stan H.